OpenStudy (anonymous):

The area of a rectangular piece of land is 240 square meters. If the length of the land was 5 meters less and the width was 3 meters more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. (5 points) Part B: What is the width of the land in meters? Show the steps of your work. (5 points)

OpenStudy (anonymous):

@sammixboo @kropot72 @CGGURUMANJUNATH @sleepyjess

OpenStudy (kropot72):

Let the length be y meters. Let the width be x meters. From the given information we can write the following: y -5 = x + 3 ............(1) @knightmare6 Do you want me to take you thru it step by step?

OpenStudy (anonymous):

yes please if you can

OpenStudy (anonymous):

helloo? @kropot72

OpenStudy (anonymous):

@confluxepic

OpenStudy (kropot72):

Rearranging equation (1) we get: y = x + 8 .................(2) The area is given by: xy = 240 ............(3) Rearranging (3) gives: \[\large x=\frac{240}{y}\ ..............(4)\] We can now substitute the expression for y in equation (2) into equation (4) giving: \[\large x=\frac{240}{x+8}\ ..........(5)\]

OpenStudy (kropot72):

Now if you cross-multiply equation (5) and rearrange you will be able to obtain a quadratic that can be solved for x.

OpenStudy (anonymous):

cross multiply what?

OpenStudy (kropot72):

Cross-multiplying (5) gives: x(x + 8) = 240 ...........(6) Now a quadratic can be formed from (6)

OpenStudy (anonymous):

x^2+8x=240

OpenStudy (kropot72):

Correct, so far. Then rearrange to give: x^2 + 8x -240 = 0 ..........(7) This is a quadratic which can be solved to find the value of x.

OpenStudy (anonymous):

so thats my answer for A?

OpenStudy (kropot72):

I believe that is sufficient for Part A, the reason being that Part B requires you to solve the quadratic.

OpenStudy (anonymous):

so i got x=-4?

OpenStudy (anonymous):

Is that correct i did -b/2a which is -8/2 and that means x=-4

OpenStudy (kropot72):

Factorizing equation (7) gives: (x + 20)(x - 12) = 0 ...........(8) Can you find the value of x now?

OpenStudy (anonymous):

what

OpenStudy (anonymous):

that foiled is x^2+8x-240

OpenStudy (kropot72):

Exactly! What values of x satisfy equation (8)?

OpenStudy (anonymous):

uhmm idk

OpenStudy (anonymous):

12 would and -20

OpenStudy (kropot72):

What is the result if we plugged the value of -20 for x in equation (8)?

OpenStudy (anonymous):

0

OpenStudy (kropot72):

Yes x = -20 or 12. A negative value for the width x is ridiculous, so what must the value of x be?

OpenStudy (anonymous):

12

OpenStudy (kropot72):

Correct. x = 12 meters.